Title

Article text.

INAF queue observing, 2016 December 7 - 8

Observer: R. Carini, A. Marchetti
Telescope Operator: S. Allanson
Support Astronomer: B. Rothberg on call
Instrument: LUCi1

Summary

LUCI2 is not available

At sunset the sky is partly cloudy

Night duration (12\xB0 el twilight) 18:10-6:00 ( 01:45-12.45 UT) = 11h 50m

Observing time: 3h 34m (30%)

Weather time loss: 8h 16m (70%)

technical time loss 0m (0%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
CALIBRATIONS Massaro MODS1-MODS2 Flats 1" Red    
Antoniucci LUCi1 PVCep zj 1200sec COMPLETED
Antoniucci LUCI1 PVCep HK 240sec COMPLETED
Antoniucci LUCI1 imaging PVCep     COMPLETED
Antoniucci LUCI1 telluric HIP111257 zJ+HK   0024 is not good, 0025 needs to be checked
CALIBRATIONS Massaro MODS1-MODS2 Flats 1.2" Red    
Ricci A2 telluric HIP41798 J   Last image taken with flexure compensation off
Ricci A2 target NGC2712 J 300sx6  
Ricci A2 target NGC2712 J 300sx6  
Ricci A2 target NGC2712 J 300sx2 missing 4 images of 300 sec for target completion

Details

00:25 we wake up MODS1 and MODS2 to take calibrations for Massaro program

00:33 send the new calibration script for 1" Massaro spectroscopic red flats for MODS1 and MODS2

  • mods1r.20161208.0002-0006
  • mods2r.20161208.0002-0006

00:42 the sky is partly cloudy, the TO opens the dome

00:49 we switch to LUCI1

LUCI ANTONIUCCI PVCEP

00:50 we send the coordinates of Antoniucci PVCep

01:04 we send the science of Antoniucci PVCep

01:09 LUCI is doing nothing. The istrument has not configured itself. The TO cancels the preset.

01:10 we send the script again. Seeing is 2". Acquisition

  • luci1.20161208.0001-0004

01:18 we send the science. Seeing is 1.5"

  • luci.20161208.0005-0008 zj filter

  • luci.20161208.0009-0012 HK filter

01:50 we send the imaging script for Antoniucci PVCep. Seeing is oscillating between 1.3" and 3"

  • luci.20161208.0013-0015

  • 01:54 guide star lost

  • 01:58 we continue the script

  • luci.20161208.0016-0018

02:00 we send the telluric for PVCep. Seeing is 1.9", but oscillating. Acquisition

  • luci.20161208.0019-0021

  • Science.
  • luci.20161208.0022-0023 zJ

*02:20 We get an error when moving to HK filter. The grating does not stabilize on 1.93 central wavelength for HK filter

02:21 we lost the guide star because of clouds

02:24 with the help of the SA we take two exposures manually. The first has an extra offset and is too faint. We take an extra one manually

  • luci.20161208.0024-0026 HK 0024 is too faint, 0025 needs to be checked

the script takes other 4 HK telluric spectra

  • ==luci.20161208.0027-0030

02:29 the sky is covered with clouds, and humidity has raised over the warning level. The TO has to close the dome

02:42 we send the calibrations for MODS1+MODS2 Massaro standard 1.2" red flats

  • mods1r.20161208.0007-0011
  • mods2r.20161208.0007-0011

08:13 the sky clears up a little bit. The TO opens the dome

LUCI RICCI TELLURIC HIP41978

08:28 we acquire the telluric of LUCI Ricci A2 HIP41798. Seeing is 0.9"

08:33 some problems with the pointing, we send the script again

08:35 the guide star is faint, we have to wait for the collimation that will take longer

08:39 we start the acquisition. Seeing is 0.9

  • luci.20161208.0031-0032

08:40 the guide star is too faint, it fails. We have to change the guide star

08:48 we take the telluric again. Acquisition

  • luci.20161208.0033-0036

Science

  • luci.20161208.0037-0038 to be checked

they have very high counts.

09:01 To be safe we take manually other frames with lower exptime. * luci.20161208.0036-0039

09:05 We take even lower exptime 4ditx2Ndit

  • luci.20161208.0040-0043 the last one is taken with flexure compensation OFF

09:11 the humidity raised too much. The TO closes the dome

10:45 sky is clearer and humidity has dropped. The TO opens the dome

LUCI RICCI A2 target

10:49 we send the coordinates of LUCI Ricci A2 target.

10:57 we run the science of Ricci A2. The guide star is a double star. We have to change the guide star.

11:00 we send the science again. Seeing is 0.8".

11:05 Acquisition

  • luci.20161208.0044-0047

11:19 LUCI is not able to stabilize the grating at the requested central wavelength of 1.28. We have to manually set the requested central wavelength. We send the script again, adjusting the offset to avoid to double it, due to the double script sending

11:21 LUCI got stuck. The SA is restarting the scheduler

11:31 we repeat the acquisition. Seeing about 2.1"

  • luci.20161208.0048-0051

11:42 we send the science. Seeing is 1.2"

  • luci.20161208.0052-0057

12:00 seeing is 0.8"

12:12 seeing is 0.9"

12:20 we send the science again (the PI requires to send 3 times the script)

  • luci.20161208.0058-0063

12:21 seeing 0.8"

12:31 seeing 0.8"

12:32 seeing peak at 1.8", suddenly back to 1.2"

12:33 seeing peak at 2.5"

12:34 seeing 0.8"-1"

12:42 seeing 0.8"-0.9"

12:49 seeing 1"

12:54 send the script again after sending a 20 pixels offset (the script does not go back to 0). Seeing 0.8"

  • luci.20161208.0064-0066

13:07 we stop the script because it is twilight time

13:09 put MODS1 and MODS2 to sleep

13:11 put blind filters in LUCI

-- %USERSIG{AlidaMarchetti - 2016-12-08}%

Comments

 
Topic revision: r6 - 19 Dec 2016, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback